# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v1

@@ 1 Literature Review
This section is dedicated to the literature review of the previous research on self-normalized inequality and empirical Bernstein inequality for marginal mean of real-valued random variables.
@@ 2.1 Martingale Difference
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(W_{i},\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$ holds for any given $i\geq 1$. Then, by denoting $\mathbf{M}_{n}=\sum_{i=1}^{n}Z_{i}$, $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z_{i}^{2}$, for any given $\delta>0$ and $n\geq 1$, we have
@@
$\mathbb{P}\Big(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\sqrt{\frac{2[\mathbb{V}]_{n}\log(\frac{1}{\delta})}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\Big)\geq 1-2\delta,\ \ \nu_{n}(\delta)=\frac{1}{1-\sqrt{2\log(\frac{1}{\delta})/n}}.$ (9)
@@ 2.2.2 Inequalities
Theorems 2 and 3 extend the martingale-difference bound of Theorem 1 to weakly dependent time series. Under mild $\phi/\tilde{\phi}-$mixing assumptions, we obtain self-normalized Bernstein-type inequalities in which the variance term is a block-based empirical long-run variance $\hat{V}_{n}$ that is fully observable from the data. The structure of the bounds is exactly analogous to the martingale case, in which the leading self-normalized term $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$ and a coverage probability $1-\delta$. In particular, we do not impose or use any explicit decay rate for the mixing coefficients. Mixing coefficient affects the high-probability bound in a very weak way. Conceptually, these results provide fixed-$n$ empirical Bernstein inequalities for the marginal mean of weakly dependent sequences, with a variance scale that automatically adapts to the appropriate long-run variance whenever it exists. This contrasts with classical Bernstein-type inequalities for mixing processes, which typically involve non-empirical long-run variance bounds and explicit mixing coefficients, or pay a loss in rate due to blocking. Here the only variance term is the simple block-based empirical long-run variance, and the resulting bounds retain a parametric $\sqrt{n}$-rate and the same sharp constant $\sqrt{2\log(\frac{1}{\delta})}$ as in the martingale-difference case. The knowledge of mixing coefficient affects the construction high-probability bound in a very indirect way. For example, under many circumstances, it is natural to assume the mixing coefficients are summable, which means $\max_{n}\Phi_{n}\lor\tilde{\Phi}_{n}<\infty$. Thus, replacing them with any user-defined diverging sequence, like $\log\log n$, yield a valid high-probability bound and the sharpness is not harmed at all since leading term is still $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$.
Unfortunately, we sometimes have to face some “strongly-dependent” situations where $\lim_{n}\sum_{k=1}^{n}\tilde{\phi}(k)<\infty$ (or $\lim_{n}\sum_{k=1}^{n}{\phi}(k)<\infty$) does NOT hold, like $\tilde{\phi}(k)=1/k$. Under this circumstance, we can not even guarantee the summability of autocovariances and the existence of the classical long-run variance. A more severe question is that it is often difficult to detect diverging speed of $\sum_{k=1}^{n}\tilde{\phi}(k)$ and $\sum_{k=1}^{n}\phi(k)$. Hence, similar to our discussion in Section 2.1, we can get out of this awkward situation by ignoring the terms containing unknown parameters. The following Theorem 4 precisely describe the cost of this ignorance. Similar to Corollary 2, the cost is usually on an exponential-level, which is often quite affordable.
@@ A.2 Proof of Results in Section 2
$L_{m,\lambda}:=\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big).$ (A.11)
Then, using (A.10) and basic properties of MDS yields
@@
Thus, we manage to prove that, for any given $\lambda\in[0,1/b]$, $\{L_{m,\lambda},\mathcal{A}_{m}\}_{m\geq 1}$ is a super martingale such that
$\mathbb{E}[L_{m,\lambda}]\leq E[L_{1,\lambda}]\leq 1,\ \forall\ m\geq 1.$ (A.12)
Step 2 (Extension of filter and randomization of $\lambda$) By defining
$L_{m}(\lambda,\cdot):=\exp\Big(\lambda\mathbf{M}_{m}(\cdot)-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}(\cdot)\Big)\ \ \text{and}\ \ g_{m}(\lambda,\omega):=\mathbb{E}[L_{m}(\lambda,\cdot)](\omega),$
@@
    $\max_{m\geq 1}\sup_{\lambda\in[0,1/b]}g_{m}(\lambda,\omega)\leq 1\$holds almost surely.
Here point (ii) is due to (A.12). Then, for any random variable $\Lambda$ taking value in $[0,1/b]$, by denoting $\mathcal{G}_{m}:=\sigma(\Lambda)\lor\mathcal{A}_{m}$ and $L_{m,\Lambda}:=L_{m}(\Lambda(\cdot),\cdot)$, $\{L_{m,\Lambda}\}_{m\geq 1}$ is adapted to filter $\{\mathcal{G}_{m}\}_{m\geq 0}$. $\mathcal{G}_{0}$ is any sigma algebra contained by the intersection of $\bigcap_{i\geq 1}\mathcal{G}_{t}$ and independent of $\sigma(\bigcup_{i\geq 1}\mathcal{G}_{i})$. This can always be done by properly enlarging the original probability space, like introducing product space. Meanwhile, Lemma A.4 implies that
@@
holds for any given $t\geq 1$. Combining law of iterated expectation and point (ii) above yields
$\mathbb{E}[L_{m,\Lambda}]=\int_{\Omega}\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)d\mathbb{P}=\int_{\Omega}g_{m}(\Omega(\omega),\omega)d\mathbb{P}\leq\max_{m\geq 1}\sup_{\lambda\in[0,1/b]}g_{m}(\lambda,\omega)\mathbb{P}(\Omega)\leq 1.$ (A.14)
@@
$\mathcal{E}=\left\{\sqrt{\frac{2\log(\frac{1}{\delta}\sum_{i=1}^{n}(Z_{i}-\mu)^{2})}{n^{2}}}<\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}.$ (A.21)
@@
A noteworthy fact is that, by denoting $\mathcal{F}_{ni}=\sigma(Z_{nj}:j\leq i)$, ② and ③ above implies the following identity holds for every given $n\geq 1$,

### v2

@@ 1 Literature Review
This section is dedicated to the literature review of the previous research on self-normalized and empirical Bernstein (EB) inequalities for marginal mean of real-valued random variables. Marginal mean here means $\mathbb{E}[Z_{1}]$ (or $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[Z_{i}]$) when samples $Z_{i}$’s have common (or different) mean(s).
@@ 2.1 Martingale Difference
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i},\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$ holds for any given $i\geq 1$. Then, by denoting $\mathbf{M}_{n}=\sum_{i=1}^{n}Z_{i}$, $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z_{i}^{2}$, for any given $\delta>0$ and $n\geq 1$, we have
@@
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\max\left\{\sqrt{\frac{2[\mathbb{V}]_{n}\log(\frac{1}{\delta})}{n^{2}}},\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}\right)\geq 1-2\delta,\ \ \nu_{n}(\delta)=\frac{1}{1-\sqrt{2\log(\frac{1}{\delta})/n}}.$ (9)
@@ 2.2.2 Inequalities
Theorems 2 and 3 extend the martingale-difference bound of Theorem 1 to weakly dependent time series. Under mild $\phi/\tilde{\phi}-$mixing assumptions, we obtain self-normalized Bernstein-type inequalities in which the variance term is a block-based empirical long-run variance $\hat{V}_{n}$ that is fully observable from the data. The structure of the bounds is exactly analogous to the martingale case, in which the leading self-normalized term $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$ and a coverage probability $1-\delta$. In particular, we do not impose or use any explicit decay rate for the mixing coefficients. Mixing coefficient affects the high-probability bound in a very weak way. Conceptually, these results provide fixed-$n$ empirical Bernstein inequalities for the marginal mean of weakly dependent sequences, with a variance scale that automatically adapts to the appropriate long-run variance whenever it exists. This contrasts with classical Bernstein-type inequalities for mixing processes, which typically involve non-empirical long-run variance bounds and explicit mixing coefficients, or pay a loss in rate due to blocking. Here the only variance term is the simple block-based empirical long-run variance, and the resulting bounds retain a parametric $\sqrt{n}$-rate and the same sharp constant $\sqrt{2\log(\frac{1}{\delta})}$ as in the martingale-difference case. The knowledge of mixing coefficient affects the construction high-probability bound in a very indirect way. For example, under many circumstances, it is natural to assume the mixing coefficients are summable (i.e. $\sum_{k=1}^{\infty}\tilde{\phi}(k)<\infty$, $\sum_{k=1}^{\infty}{\phi}(k)<\infty$). Thus, replacing them with any user-defined diverging sequence, like $\log\log n$, yield a valid high-probability bound and the sharpness is not harmed at all since leading term is still $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$.
Unfortunately, we sometimes have to face some “strongly-dependent” situations where the assumption of summability of mixing coefficients may NOT hold, like the case where we can only assume $\tilde{\phi}(k)\lesssim 1/k$. Under this circumstance, we can not even guarantee the summability of autocovariances and the existence of the classical long-run variance. A more severe question is that it is often difficult to detect diverging speed of $\sum_{k=1}^{n}\tilde{\phi}(k)=:\tilde{\Phi}_{n}$ and $\sum_{k=1}^{n}\phi(k)=:\Phi_{n}$. Hence, according to our discussion in Section 2.1, we can get out of this awkward situation by ignoring the terms containing unknown parameters. The following Theorem 4 precisely describe the cost of this ignorance. Similar to Corollary 2, the cost is also on an exponential-level, which is usually affordable.
@@ A.2 Proof of Results in Section 2
$L_{m,\lambda}:=\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big)$ (A.11)
Then, using (A.10) and basic properties of MDS yields that, for any given $m\geq 1$,
@@
Thus, we manage to prove that, for any given $\lambda\in[0,1/b]$, $\{L_{m,\lambda},\mathcal{A}_{m}\}_{m\geq 1}$ is a super martingale Step 2 (Extension of filter and randomization of $\lambda$) Define
$L_{m}(\lambda,\cdot):=\exp\Big(\lambda\mathbf{M}_{m}(\cdot)-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}(\cdot)\Big)\ \ \text{and}\ \ g_{m}(\lambda,\omega):=\mathbb{E}[L_{m}(\lambda,\cdot)|\mathcal{A}_{m-1}](\omega),$
@@
    For each given $\lambda\in(0,1/b]$, $\mathbb{P}(\omega\in\{g_{m}(\lambda,\omega)\leq L_{m-1}(\lambda,\omega)\})=1$.
Then, for any random variable $\Lambda$ taking value in $[0,1/b]$, by denoting $\mathcal{G}_{m}:=\sigma(\Lambda)\lor\mathcal{A}_{m}$ and $L_{m,\Lambda}(\cdot):=L_{m}(\Lambda(\cdot),\cdot)$, $\{L_{m,\Lambda}\}_{m\geq 1}$ is adapted to filter $\{\mathcal{G}_{m}\}_{m\geq 0}$. $\mathcal{G}_{0}$ is any sigma algebra contained by the intersection of $\bigcap_{i\geq 1}\mathcal{G}_{t}$ and independent of $\sigma(\bigcup_{i\geq 1}\mathcal{G}_{i})$. This can always be done by properly enlarging the original probability space, like introducing product space. Meanwhile, Lemma A.4 implies that
@@
holds for any given $m\geq 1$. Together with Point (ii) above, we thus obtain
$\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)=g_{m}(\Lambda(\omega),\omega)\leq L_{m-1}(\Lambda(\omega),\omega)=:L_{m-1,\Lambda}(\omega),$
which indicates that $(L_{m,\Lambda},\mathcal{G}_{m})_{m\geq 1}$ is indeed an non-negative super-martingale. Then, combining law of iterated expectation above yields
$\max_{m}\mathbb{E}[L_{m,\Lambda}]\leq 1.$ (A.13)
@@
$\mathcal{E}=\left\{\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}<\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}.$ (A.20)
@@
A noteworthy fact is that, by denoting $\mathcal{F}_{ni}=\sigma(Z_{nj}:j\leq i)$, Lemma A.5 and ②, ③ above imply the following identity holds for every given $n\geq 1$,